# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ D.1 Formal Statement of Theorem 7
This is the result of substituting the terms of (111)-(115) and Theorem 2 to Theorem 9, and relating the span semi-norm to the $\|\cdot\|_{\infty,\overline{E}}$ semi-norm by (49). ∎
D.2 Proof of Theorem 7
We use the result from Theorem 10, to set $\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathrm{sp}}\Bigr{]}\leq\epsilon^{2}$. For contamination uncertainty sets we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, conbining with Theorem 4, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
D.3 Formal Statement of Theorem 8
@@
$=2{\|V_{T}-V^{*}\|}_{\mathrm{sp}}$ (124)
Thus, the following recursion can be formed
@@
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(2{\|V_{T}-V^{*}\|}_{\mathrm{sp}}+\left|\nu_{t}\right|)$ (125)
@@
$\mathbb{E}\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]$ $\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (126)
@@
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]$ $\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(2\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (127)
@@
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\mathrm{sp}}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (128)
@@
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$ (129)
$=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right)$ (130)
@@
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$
$+\frac{4\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$ (131)
@@
$\mathbb{E}\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}$ $\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{4K\sqrt{c_{u}}\log T}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{8\sqrt{A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}\sqrt{T+K}}$
$+\frac{4\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{\sqrt{\alpha_{2}}}+48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$ (133)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{Cont}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}$ (135)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{TV}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}}$ (136)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\mathrm{sp}}\Bigr{]}\leq\frac{K\sqrt{c_{u}}}{(T+K)\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\mathrm{sp}}+\frac{2\sqrt{A^{\mathrm{Wass}}\alpha_{4}c_{u}}}{\alpha_{2}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}}$ (137)
@@ D.4 Proof of Theorem 8
Appendix E Some Auxiliary Lemmas for the Proofs
Lemma 4 (Lemma 6 in (Zhang et al., 2021)).
Under the setup and notation in Appendix B.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$, we have

### v2

@@ F.1 Formal Statement of Theorem 6
By Lemma 9 and (120), we have that for any value function $V$ its $\|\cdot\|_{\mathcal{P}}$ norm is bounded as follows:
$\|V\|_{\mathcal{P}}\leq 4C_{\mathcal{P}}t_{\rm mix}$ (137)
Substituting the terms of (127)-(F.1), (137), and Theorem 2 to Theorem 9, we would have for contamination uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{8A^{\mathrm{Cont}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T^{2}}+\frac{t^{2}_{\mathrm{mix}}}{T(1-\gamma)^{2}}\right)$ (138)
for TV distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{8A^{\mathrm{TV}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}}$ (139)
for Wasserstein distance uncertainty sets,
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\mathcal{P}}\Bigr{]}$ $\leq\frac{K^{2}c_{u}}{(T+K)^{2}c_{l}}\|V_{0}-V^{*}\|^{2}_{\mathcal{P}}+\frac{4A^{\mathrm{Wass}}\alpha_{4}c_{u}}{(T+K)\alpha_{2}^{2}}+\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}}$ (141)
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}}$ is defined in Lemma 9. We now translate the result back to the standard $l_{\infty}$ norm by applying Lemma 9 again to thee above, we obtain the desired results. ∎
F.2 Proof of Theorem 6
We use the result from Theorem 10, to set $\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 3, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
[…]
F.3 Formal Statement of Theorem 7
@@
$=2\|V_{T}-V^{*}\|_{\mathrm{sp}}$
$\leq 4\|V_{T}-V^{*}\|_{\infty}$ (146)
Where the last inequality is by Lemma 9. Thus, the following recursion can be formed
@@
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}(4\|V_{T}-V^{*}\|_{\infty}+\left|\nu_{t}\right|)$ (147)
@@
$\mathbb{E}$ $\left[\left|g_{t+1}-g^{\pi}_{\mathcal{P}}\right|\right]\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\mathbb{E}[\left|\nu_{t}\right|]\right)$
$\leq(1-\beta_{t})\left|g_{t}-g^{\pi}_{\mathcal{P}}\right|+\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$ (148)
@@
$\mathbb{E}$ $\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|$
$+\zeta_{T}\sum_{t=0}^{T-1}(\frac{1}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}(\frac{\zeta_{T}}{\zeta_{t+1}})\beta_{t}\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)$
$\leq\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\sum_{t=0}^{T-1}\beta_{t}{\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)}$
$=\zeta_{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(4\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\sum_{t=0}^{T-1}\beta_{t}$ (149)
@@
$\mathbb{E}\left[\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\right]\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\left(8\mathbb{E}\left[{\|V_{T}-V^{*}\|}_{\infty}\right]+2\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})-{\sigma}_{\mathcal{P}^{a}_{s}}(V_{T})\right]\right|\right)\log T$ (150)
@@
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}=\operatorname*{\mathcal{O}}\left(\frac{1}{T}+\frac{\log T}{T}+\frac{t_{\mathrm{mix}}\log T}{\sqrt{T}(1-\gamma)}\right)$ (151)
@@
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48(1+\frac{1}{\delta})2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
@@
$\mathbb{E}$ $\Bigl{[}\left|g_{T}-g^{\pi}_{\mathcal{P}}\right|\Bigr{]}\leq\frac{1}{T}\left|g_{0}-g^{\pi}_{\mathcal{P}}\right|+\frac{16KC_{\mathcal{P}}\sqrt{c_{u}}\log T}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}$
$+\frac{16\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}\log T}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\frac{8\sqrt{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}\log T}{c_{\mathcal{P}}\sqrt{\alpha_{2}}}+48\cdot 2^{-\frac{N_{\max}}{2}}t_{\mathrm{mix}}\log T$
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{Cont}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}$ (154)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{TV}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{TV}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}}$ (155)
@@
$\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|_{\infty}\Bigr{]}\leq\frac{2KC_{\mathcal{P}}\sqrt{c_{u}}}{(T+K)c_{\mathcal{P}}\sqrt{c_{l}}}\|V_{0}-V^{*}\|_{\infty}+\frac{2\sqrt{2A^{\mathrm{Wass}}\alpha_{4}c_{u}}}{\alpha_{2}c_{\mathcal{P}}\sqrt{T+K}}+\sqrt{\frac{c_{u}C_{3}C_{2}\varepsilon^{\mathrm{Wass}}_{\mathrm{bias}}}{\alpha_{2}c_{\mathcal{P}}^{2}}}$ (156)
Regarding the bound for the absolute bias of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}$, from Lemma 12, we have for contamination uncertainty,
$\left|\mathbb{E}\left[\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)-{\sigma}_{\mathcal{P}^{a}_{s}}(V)\right]\right|=0$ (157)
@@ F.4 Proof of Theorem 7
Appendix G Some Auxiliary Lemmas for the Proofs
Lemma 10 (Theorem IV in Berger and Wang (1992)).
Let $\mathcal{Q}$ be a bounded set of square matrix such that $\rho(Q)<\infty$ for all $Q\in\mathcal{Q}$ where $\rho(\cdot)$ denotes the spectral radius. Then we have
$\sup_{Q\in\mathcal{Q}}\rho(Q)=\lim_{k\rightarrow\infty}\sup_{Q_{i}\in\mathcal{Q}}\rho(Q_{k}\ldots Q_{1})^{\frac{1}{k}}$ (160)
Lemma 11 (Lemma 6 in Zhang et al. (2021)).
Under the setup and notation in Appendix D.1.1, if assuming the noise has bounded variance of $\mathbb{E}[\,\|w^{t}\|_{\infty,\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\infty,\overline{E}}^{2}$, we have